OpenStudy (anonymous):

solve on the interval [0,2pi) 1-cos theta=(2- sqrt 3)/2

OpenStudy (mertsj):

\[1-\cos \theta=\frac{2-\sqrt{3}}{2}\] \[2-2\cos \theta=2-\sqrt{3}\] \[-2\cos \theta=-\sqrt{3}\] \[\cos \theta=\frac{\sqrt{3}}{2}\] \[\theta=\frac{\pi}{6}, \frac{11\pi}{6}\]